The circumferences of two circles are 3 m. and 9p m. What is the ratio of the areas of the circle?


1
Expert's answer
2021-01-21T00:21:16-0500

Circumferences: "C_1=2\pi r_1,\;C_2=2\pi r_2."

"\frac{r_2}{r_1}=\frac{C_2}{C_1}=\frac{9}{3}=3."

Areas: "A_1=\pi r_1^2,\;A_2=\pi r_2^2."

"\frac{A_2}{A_1}=\frac{r_2^2}{r_1^2}=3^2=9."